OpenStudy (anonymous):

1/3 x + 2 = 10 for x. A. 36 B. 30 C. 24 D. 12

OpenStudy (anonymous):

add -2 to both sides :)

OpenStudy (anonymous):

1/3 * x=10-2 1/3 *x=8 x/3=8 x=24

OpenStudy (anonymous):

@midnight830

OpenStudy (austinl):

Subtract 2 from both sides, then multiply by 3. That should get you your answer.

OpenStudy (anonymous):

option C

OpenStudy (anonymous):

@midnight830 , got it?

OpenStudy (anonymous):

got it, @midnight830